What is next in this sequence and how do I solve probems like this, is there a formula?
1 5 14 30 55 91 140

Found 3 solutions by scott8148, rfer, checkley77:
Answer by scott8148(6628)

look at the DIFFERENCE between consecutive terms

4 , 9 , 16 , 25 , 36 , 49 ___ see a pattern?

the next term should be ___ 140 + 64


if the first set of differences don't show anything , take the differences of the differences
___ in this case ___ 5 , 7 , 9 , 11 , 13

the set of differences that shows a CONSTANT change indicates the "order" of the relation
___ in this case ___ the 2nd set has a constant change , so it is a second order (squared) relation

Answer by rfer(16322)

204 is next
-------------
1 5 14 30 55 91 140 204
4 9 16 25 36 49 64
5 7 9 11 13 15
2 2 2 2 2

Answer by checkley77(12844)

1 5 14 30 55 91 140
Calculate the difference between each value.
4 9 16 25 36 49 these are the squares of 2,3,4,5,6 & 7
Thus the next number will be 140+8^2=140+64=204.